The exterior angle of a regular polygon is one third of its interior angle.How many sides does the polygon have?
Answer by Alan3354(69443)

The exterior angle of a regular polygon is one third of its interior angle.How many sides does the polygon have?
-----------
E + I = 180
I = 3E
E + 3E = 180
E = 45º
Interior angle = 135º
---------
Interior angle = (n-2)*180/n
135n = 180n - 360
-45n = -360
n = 8